40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 12 03:08:23 2021, Koji, Summary, SUS, Coil Rs & Ls for PRM/BS/SRM P_20210311_224651.jpgP_20210311_225359.jpg
    Reply  Fri Mar 12 12:32:54 2021, gautam, Summary, SUS, Coil Rs & Ls for PRM/BS/SRM BS_actuator.pdfPRMact.pdf
       Reply  Fri Mar 12 13:01:43 2021, rana, Summary, SUS, Coil Rs & Ls for PRM/BS/SRM 
          Reply  Fri Mar 12 13:48:53 2021, gautam, Summary, SUS, Coil Rs & Ls for PRM/BS/SRM 
    Reply  Thu Mar 25 16:02:15 2021, gautam, Summary, SUS, Repeated measurement of coil Rs & Ls for PRM/BS 
Message ID: 15966     Entry time: Thu Mar 25 16:02:15 2021     In reply to: 15907
Author: gautam 
Type: Summary 
Category: SUS 
Subject: Repeated measurement of coil Rs & Ls for PRM/BS 

Method

Since I am mainly concerned with the actuator part of the OSEM, I chose to do this measurement at the output cables for the coil drivers in 1X4. See schematic for pin-mapping. There are several parts in between my measurement point and the actual coils but I figured it's a good check to figure out if measurements made from this point yield sensible results. The slow bias voltages were ramped off under damping (to avoid un-necessarily kicking the optics when disconnecting cables) and then the suspension watchdogs were shutdown for the duration of the measurement.

I used an LCR meter to measure R and L - as prescribed by Koji, the probe leads were shorted and the readback nulled to return 0. Then for R, I corroborated the values measured with the LCR meter against a Fluke DMM (they turned out to be within +/- 0.5 ohms of the value reported by the BK Precision LCR meter which I think is reasonable).

Result

                   PRM
Pin1-9 (UL)   / R = 30.6Ω / L=3.23mH  
Pin2-10 (LL)  / R = 30.3Ω / L=3.24mH
Pin3-11 (UR) / R = 30.6Ω / L=3.25mH
Pin4-12 (LR) / R = 31.8Ω / L=3.22mH
Pin5-13 (SD) / R = 30.0Ω / L=3.25mH

                       BS
Pin1-9 (UL)   / R = 31.7Ω / L=3.29mH
Pin2-10 (LL)  / R = 29.7Ω / L=3.26mH
Pin3-11 (UR) / R = 29.8Ω / L=3.30mH
Pin4-12 (LR) / R = 29.7Ω / L=3.27mH
Pin5-13 (SD) / R = 29.0Ω / L=3.24mH

Conclusions

On the basis of this measurement, I see no problems with the OSEM actuators - the wire resistances to the flange seem comparable to the nominal OSEM resistance of ~13 ohms, but this isn't outrageous I guess. But I don't know how to reconcile this with Koji's measurement at the flange - I guess I can't definitively rule out the wire resistance being 30 ohms and the OSEMs being ~1 ohm as Koji measured. How to reconcile this with the funky PRM actuator measurement? Possibilities, the way I see it, are:

  1. Magnets on PRM are weird in some way. Note that the free-swinging measurement for the PRM showed some unexpected features.
  2. The imbalance is coming from one of the drive chain - could be a busted current buffer for example.
  3. The measurement technique was wrong.
ELOG V3.1.3-